# Estimation of Tsallis entropy and its applications to goodness-of-fit tests

### v1

@@ VIII-D Data Analyses
TABLE XII: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(9,09)R=(9,0*9).
@@
TABLE XIII: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(09,9)R=(0*9,9).
@@
TABLE XIV: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(5,08,4)R=(5,0*8,4).
@@
From data set 2 of Grubbs et al. (1971), we generate three PC-II censored data set and apply the proposed exponentiality test to these three data. The data dets along with censoring schemes are provided in Tables XII-XIV, respectively. We calculate the test statistics and corresponding p-values for these three data.
@@
  • For the 3rd data, $\hat{T_{\alpha}}(F_{r,n},F)$ = 41095 and p-value is 0.323, suggesting acceptance of exponentiality hypothesis.
IX Concluding remarks

### v2

@@ 8.4 Data Analyses
Table 12: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(9,09)R=(9,0*9).
@@
Table 13: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(09,9)R=(0*9,9).
@@
Table 14: Progressive censored sample generated from data set 2 with ($n,m$) = (19, 10) and R=(5,08,4)R=(5,0*8,4).
@@
From data set 2 of Grubbs et al. (1971), we generate three PC-II censored data set and apply the proposed exponentiality test to these three data. The data dets along with censoring schemes are provided in Tables 12-14, respectively. We calculate the test statistics and corresponding p-values for these three data sets.
@@
  • For the 3rd data, $\hat{T_{\alpha}}(F_{r,n},F)$ = 0.41095 and p-value is 0.323, suggesting acceptance of exponentiality hypothesis.
9 Concluding remarks